# Transfer Learning for High-dimensional Quantile Regression with Distribution Shift

### v1

@@ 3 Transferable Set Detection
Since the oracle transferable set $\mathcal{C}_{h}$ may be unknown in practice, here we propose a detection procedure to screen out the source studies satisfying the requirements in (3). We can first get an initial estimate $\widetilde{\boldsymbol{\beta}}$ through $\ell_{1}$-QR on the target study, then remove the part accounted by $\widetilde{\boldsymbol{\beta}}$ on each source study and conduct $\ell_{1}$-QR to obtain the contrast estimators $\{\widetilde{\boldsymbol{\delta}}^{(k)}\}_{k=1}^{K}$. With contrast estimators $\{\widetilde{\boldsymbol{\delta}}^{(k)}\}_{k=1}^{K}$, screening studies with small $\|\boldsymbol{\delta}^{(k)}\|_{1}$ is relatively simple. We can directly calculate each $\|\widetilde{\boldsymbol{\delta}}^{(k)}\|_{1}$ and set a threshold to reject the dissimilar studies. As implied by Corollary 1, $h_{1}$ should be no more than the single-task error, here we set the threshold as $t_{1}\sqrt{\log p/n_{0}}$, where $t_{1}$ is a thresholding parameter. The first screening set is then defined as $\widehat{\mathcal{C}}_{1}=\{k:\|\widetilde{\boldsymbol{\delta}}^{(k)}\|_{1}\leq t_{1}\sqrt{\log p/n_{0}}\}$. We will show that under some regularity conditions, appropriate choice of $t_{1}$ will lead to consistent selection conforming with the requirement $\|\boldsymbol{\delta}^{(k)}\|_{1}\leq h_{1}$.
@@
The second requirement of the oracle transferable set $\mathcal{C}_{h}$ is to reject source studies with tiny expectation of $\tau$-th density $\mathbb{E}[f_{i}^{(k)}]$ compared to the target $\mathbb{E}[f_{i}^{(0)}]$. We employ the approxmation technique with the Powell bandwidth $b_{k}$ (Koenker,, 2005) to estimate each $\mathbb{E}[f_{i}^{(k)}]$. Specifically, for each study we calculate
@@
$\widetilde{f}^{(k)}(\widetilde{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)})=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\mathbb{I}\left(|y_{i}^{(k)}-(\boldsymbol{x}_{i}^{(k)})^{\top}(\widetilde{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)})|\leq b_{k}\right)/2b_{k},$
where the practical choice of bandwidth $b_{k}$ will be recommended in the simulation section. Denote the second screening set $\widehat{\mathcal{C}}_{2}=\{k:\widetilde{f}^{(k)}(\widetilde{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)})\geq t_{2}\widetilde{f}^{(0)}(\widetilde{\boldsymbol{\beta}})\}$ with the thresholding parameter $t_{2}$. We will show that under some regularity conditions, appropriate choice of $t_{2}$ will lead to consistent selection conforming with the requirement $\mathbb{E}[f^{(0)}_{i}]/\mathbb{E}[f^{(k)}_{i}]\leq h_{2}$. The final detected transferable set is $\widehat{\mathcal{C}}=\widehat{\mathcal{C}}_{1}\cap\widehat{\mathcal{C}}_{2}$. Note that the calculation of estimators $\widetilde{\boldsymbol{\beta}},\{\widetilde{\boldsymbol{\delta}}^{(k)}\}_{k=1}^{K}$ is exactly the first round of Algorithm 1 for all source studies. Also in our simulations, Algorithm 1 usually converges in one step, hence here we propose a unified one-shot algorithm for transferable set detection and parameter prediction, which is shown in Algorithm 2.
@@
Input: Target data $\{((\boldsymbol{x}_{i}^{(0)})^{\top},y_{i}^{(0)})_{i=1}^{n_{0}}\}$, source data $\{((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)})_{i=1}^{n_{k}}\}_{k=1}^{K}$, tuning parameters $\lambda_{\boldsymbol{\beta}}$ and $\{\lambda_{k}\}_{0\leq k\leq K}$, bandwidth parameters $\{b_{k}\}_{0\leq k\leq K}$.
@@
    Compute each $\widetilde{\boldsymbol{\delta}}^{(k)}\leftarrow\ell_{1}\text{-QR }\text{ on }[((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)}-(\boldsymbol{x}_{i}^{(k)})^{\top}\widetilde{\boldsymbol{\beta}})_{i=1}^{n_{k}},\lambda_{k}],1\leq k\leq K$.
@@
    Compute $\widehat{\mathcal{C}}_{1}=\{k:\|\widetilde{\boldsymbol{\delta}}^{(k)}\|_{1}\leq t_{1}\sqrt{\log p/n_{0}}\}$ and $\widehat{\mathcal{C}}_{2}=\{k:\widetilde{f}^{(k)}(\widetilde{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)})\geq t_{2}\widetilde{f}^{(0)}(\widetilde{\boldsymbol{\beta}})\}$, where $\widetilde{f}^{(k)}(\widetilde{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)})=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\mathbb{I}(|y_{i}^{(k)}-(\boldsymbol{x}_{i}^{(k)})^{\top}(\widetilde{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)})|\leq b_{k})/2b_{k}$.
@@
    Determine the transferable set $\widehat{\mathcal{C}}=\widehat{\mathcal{C}}_{1}\cap\widehat{\mathcal{C}}_{2}$, then solve the pooled $\ell_{1}$-QR problem
@@
We next investigate the theoretical properties of Algorithm 2, which determines the transferable set $\widehat{\mathcal{C}}$. The key ingredient is to establish the detection consistency of $\widehat{\mathcal{C}}$ towards the oracle but unknown transferable set $\mathcal{C}_{h}$.
@@
For each source study $k$, let $\Omega_{1}^{(k)}:=C_{3}\frac{s_{k}}{\mathbb{E}[f_{i}^{(k)}]}\sqrt{\frac{\log p}{n_{k}\wedge n_{0}}}+2\eta_{k}$, where $\eta_{k}$ is defined in Condition 4, and $C_{3}$ is a positive constant specified in the supplementary materials. Denote the set $\mathcal{C}_{h_{1}}:=\{k\in\{1,\dots,K\}:\|\boldsymbol{\delta}^{(k)}\|\leq h_{1}\}$. $\mathcal{C}_{h_{1}}$ satisfies that
@@
where $r_{k}$ is defined in Theorem 1 ($r_{0}=0$), and $C_{4}$ is a positive constant specified in the supplementary materials. Denote the set $\mathcal{C}_{h_{2}}:=\{k\in\{1,\dots,K\}:\mathbb{E}[f_{i}^{(0)}]/\mathbb{E}[f_{i}^{(k)}]\leq h_{2}\}$. $\mathcal{C}_{h_{2}}$ satisfies that
$\frac{\inf_{k\in\mathcal{C}_{h_{2}}}(\mathbb{E}[f_{i}^{(k)}]-\Omega_{2}^{(k)})}{\mathbb{E}[f_{i}^{(0)}]+\Omega_{2}^{(0)}}>\frac{\sup_{k\in\mathcal{C}_{h_{2}}^{c}}(\mathbb{E}[f_{i}^{(k)}]+\Omega_{2}^{(k)})}{\mathbb{E}[f_{i}^{(0)}]-\Omega_{2}^{(0)}}.$
@@
Condition 5 and 6 assume a significant gap of quantile coefficients and $\tau$-th densities to ensure the identifiability of transferable source studies satisfying the oracle requirement in (3). In fact, $\Omega_{1}^{(k)},\Omega_{2}^{(k)}$ are estimation error bounds for $\|\widetilde{\boldsymbol{\delta}}^{(k)}-\boldsymbol{\delta}^{(k)}\|_{1}$ and $|\widetilde{f}^{(k)}-\mathbb{E}[f_{i}^{(k)}]|$ respectively. Similar conditions are also assumed in Tian and Feng, (2023); Zhang and Zhu, (2023); Huang et al., (2023). Moreover, since we don’t adopt the detection method by comparing loss functions, we relax the restrictions on the similarity of covariance matrices (for example Assumption 6 in Huang et al., (2023)).
@@
Assuming Conditions 1-6 and assumptions in Theorem 1, choose the thresholding parameters $t_{1},t_{2}$ in Algorithm 2 as
@@
$\frac{\inf_{k\in\mathcal{C}_{h_{2}}}(\mathbb{E}[f_{i}^{(k)}]-\Omega_{2}^{(k)})}{\mathbb{E}[f_{i}^{(0)}]+\Omega_{2}^{(0)}}$ $\geq t_{2}\geq\frac{\sup_{k\in\mathcal{C}_{h_{2}}^{c}}(\mathbb{E}[f_{i}^{(k)}]+\Omega_{2}^{(k)})}{\mathbb{E}[f_{i}^{(0)}]-\Omega_{2}^{(0)}}.$
Then for the detected transferable set $\widehat{\mathcal{C}}$ in Algorithm 2, with any constant $\varepsilon>0$,
$\mathbb{P}(\widehat{\mathcal{C}}=\mathcal{C}_{h})\geq 1-\varepsilon-2(K+1)A_{1}\exp(-A_{2}\log p).$
Consequently, our practical $\widehat{\boldsymbol{\beta}}$ obtained from Algorithm 2 can enjoy the same $\ell_{1}/\ell_{2}$-convergence rates in Corollary 1.
@@ 4.1 Debiased Estimator based on Neyman Orthogonality
We now turn to the statistical inference for the target parameter, $\boldsymbol{\beta^{*}}$, focusing specifically on the first component, $\beta^{*}_{1}$, without loss of generality. Previous studies (Tian and Feng,, 2023; Li et al.,, 2024; Huang et al.,, 2023) have addressed this inference task building on the principles of debiased lasso (Zhang and Zhang,, 2014; van de Geer et al.,, 2014; Javanmard and Montanari,, 2014). In these approaches, the primary method involves estimating the precision matrix of the loss function and applying a one-step correction to the $\ell_{1}$-penalized estimate. Unfortunately, these methods only achieve $\sqrt{n_{0}}$-normality for the debiased estimators, offering little improvement over debiasing based solely on the target study. The key reason for this limitation is that their debiasing procedures need the covariance matrix to be identical for all samples, a condition that is violated when covariate shift occurs. Also, they do not offer any measures to adjust for the parameter contrast $\boldsymbol{\delta}^{(k)}$. As a result, the $\sqrt{n_{0}}$-normality is suboptimal, especially given the wealth of transferable information from source studies. In an ideal scenario where $\boldsymbol{\delta}^{(k)}=0$ and the residuals $\epsilon_{i}^{(k)}$ follow a normal distribution $\mathcal{N}(0,1)$, meaning all source studies are as informative as the target study, we should attain a $\sqrt{N}$-normality by pooling all studies together. The challenge, then, is how to effectively leverage the information from these transferable source studies to enhance inference efficiency, particularly under distributional shifts.
To address this challenge, we propose a novel approach motivated by the Neyman orthogonal idea (Belloni et al.,, 2019). Before introducing the whole algorithm, we use the $k$-th source study as a methodological warm-up. Denote $\boldsymbol{v}_{-1}$ as the last $p-1$ dimensions of the vector $\boldsymbol{v}\in\mathbb{R}^{p}$. We can rewrite the original model (1) as a partially linear quantile regression
@@
The high-dimensional nuisance parameter $\boldsymbol{w}_{-1}^{(k)}$ will bring trouble for identifying $\beta^{*}_{1}$, since the first-order derivative of (9) with respect to this nuisance parameter at $\alpha=\beta^{*}_{1}$ is non-zero:
$\left.\partial_{\boldsymbol{w}}\mathbb{E}\sum_{i=1}^{n_{k}}\left[\mathbb{I}\left\{y_{i}^{(k)}\leq x_{i,1}^{(k)}\beta^{*}_{1}+(\boldsymbol{x}_{i,-1}^{(k)})^{\top}\boldsymbol{w}+r_{i}^{(k)}\right\}-\tau\right]x_{i,1}^{(k)}\right|_{\boldsymbol{w}=\boldsymbol{w}_{-1}^{(k)}}=\sum_{i=1}^{n_{k}}f_{i}^{(k)}x_{i,1}^{(k)}\boldsymbol{x}_{i,-1}^{(k)}\neq 0.$
If conducting valid inference on $\beta^{*}_{1}$ directly with moment condition (9), the estimator for $\boldsymbol{w}_{-1}^{(k)}$ needs to converge faster than $1/\sqrt{n_{k}}$, which is difficult to achieve in theory. To address this problem, we adopt the Neyman orthogonality to make the moment condition immune to the first-order error from the nuisance estimator. The technique is based on the linear projection of the regressor of interest $x_{i,1}^{(k)}$ on the $\boldsymbol{x}^{(k)}_{i,-1}$ weighted by the $\tau$-th density $f_{i}^{(k)}$:
@@
$\psi_{i}^{(k)}(\alpha,\boldsymbol{w}_{-1}^{(k)},\boldsymbol{\theta}_{0}^{(k)},f_{i}^{(k)}):=\left[\mathbb{I}\{y_{i}^{(k)}\leq x_{i,1}^{(k)}\alpha+(\boldsymbol{x}_{i,-1}^{(k)})^{\top}\boldsymbol{w}_{-1}^{(k)}+r_{i}^{(k)}\}-\tau\right]\left[f_{i}^{(k)}x_{i,1}^{(k)}-f_{i}^{(k)}(\boldsymbol{x}^{(k)}_{i,-1})^{\top}\boldsymbol{\theta}_{0}^{(k)}\right],$ (11)
@@
and solve for $\mathbb{E}\sum_{i=1}^{n_{k}}[\psi_{i}^{(k)}(\alpha,\boldsymbol{w}_{-1}^{(k)},\boldsymbol{\theta}_{0}^{(k)}),f_{i}^{(k)}]=0$, where this moment condition can satisfy the orthogonality with respect to high-dimensional nuisance parameters $\boldsymbol{w}_{-1}^{(k)}$ and $\boldsymbol{\theta}_{0}^{(k)}$, i.e.
$\left.\partial_{\boldsymbol{w}}\mathbb{E}\sum_{i=1}^{n_{k}}\left[\psi_{i}^{(k)}(\beta_{1}^{*},\boldsymbol{w},\boldsymbol{\theta}_{0}^{(k)},f_{i}^{(k)})\right]\right|_{\boldsymbol{w}=\boldsymbol{w}_{-1}^{(k)}}=\sum_{i=1}^{n_{k}}f_{i}^{(k)}\boldsymbol{x}_{i,-1}^{(k)}v_{0i}^{(k)}=0,$
$\left.\partial_{\boldsymbol{\theta}}\mathbb{E}\sum_{i=1}^{n_{k}}\left[\psi_{i}^{(k)}(\beta_{1}^{*},\boldsymbol{w}_{-1}^{(k)},\boldsymbol{\theta},f_{i}^{(k)})\right]\right|_{\boldsymbol{\theta}=\boldsymbol{\theta}_{0}^{(k)}}=\left.\partial_{\boldsymbol{\theta}}\mathbb{E}_{\boldsymbol{x}_{i}^{(k)}}[~{}0~{}]\right|_{\boldsymbol{\theta}=\boldsymbol{\theta}_{0}^{(k)}}=0.$
Consequently, the estimators for $\boldsymbol{w}_{-1}^{(k)}$ and $\boldsymbol{\theta}_{0}^{(k)}$ will only introduce a second-order bias, allowing the use of high-dimensional sparse estimates. It is important to note that none of the debiased approaches can be applied to the approximate error term, $r_{i}^{(k)}=x_{i,1}^{(k)}\delta^{(k)}_{1}$, because each $\delta^{(k)}_{1}$ shares the same covariate $x_{i,1}^{(k)}$ with the target parameter $\beta^{*}_{1}$. We will address the requirements and potential solutions for handling this term in the context of knowledge transfer later.
Now let’s discuss why this orthogonal approach can conduct knowledge transfer for statistical inference in the presence of distribution shift. The key point is that this approach can avoid the computation of precision matrices and free the identical covariance matrix assumption across studies, which exactly accommodates the circumstance under covariate shift. To tackle the parameter shift and residual shift, we only incorporate moment conditions of target and sources in the oracle transferable set $\mathcal{C}_{h}$, where the oracle moment condition is to solve
$\mathbb{E}\sum_{k\in\{0\}\cup\mathcal{C}_{h}}\sum_{i=1}^{n_{k}}\left[\psi_{i}^{(k)}(\alpha,\boldsymbol{w}_{-1}^{(k)},\boldsymbol{\theta}_{0}^{(k)},f_{i}^{(k)})\right]=0.$ (12)
The parameter identified from (12) leverages information from informative source studies rather than only debiasing on the target, hence could achieve better inference performance than previous works.
We outline several points here for practical implementation. Firstly, we estimate the $\tau$-th density $f_{i}^{(k)}$ by
$\widehat{f}_{i}^{(k)}=\frac{2h_{k}}{(\boldsymbol{x}_{i}^{(k)})^{\top}\widehat{\boldsymbol{\beta}}^{(k)}_{\tau+h_{k}}-(\boldsymbol{x}_{i}^{(k)})^{\top}\widehat{\boldsymbol{\beta}}^{(k)}_{\tau-h_{k}}}\quad,$ (13)
@@
where $\widehat{\boldsymbol{\beta}}^{(k)}_{\tau\pm h_{k}}$ are computed via $\ell_{1}$-QR at $(\tau\pm h_{k})$-th quantile with tuning parameters $\lambda_{\tau}^{(k)}$ on the $k$-th study respectively. The bandwidth parameter, $h_{k}$, is set to $\min\{n_{k}^{-1/6},\tau(1-\tau)/2\}$, as suggested by Koenker, (2005); Giessing and Wang, (2023). Next, each high-dimensional projection parameter $\boldsymbol{\theta}_{0}^{(k)}$ is determined by $\ell_{1}$-penalized weighted least squares
$\widehat{\boldsymbol{\theta}}^{(k)}\in\arg\min_{\boldsymbol{\theta}}\sum_{i=1}^{n_{k}}\left[\widehat{f}^{(k)}_{i}x_{i,1}^{(k)}-\widehat{f}^{(k)}_{i}(\boldsymbol{x}^{(k)}_{i,-1})^{\top}\boldsymbol{\theta}\right]^{2}+\lambda_{\boldsymbol{\theta}}^{(k)}\|\boldsymbol{\theta}\|_{1},$ (14)

### v2

@@ 3 Transferable Set Detection
Since the oracle transferable set $\mathcal{C}_{h}$ may be unknown in practice, here we propose a detection procedure to screen out the source studies satisfying the requirements in (2). We can first get an initial estimate $\widetilde{\boldsymbol{\beta}}$ through $\ell_{1}$-QR on the target study, then remove the part accounted by $\widetilde{\boldsymbol{\beta}}$ on each source study and conduct $\ell_{1}$-QR to obtain the contrast estimators $\{\widetilde{\boldsymbol{\delta}}^{(k)}\}_{k=1}^{K}$. With contrast estimators $\{\widetilde{\boldsymbol{\delta}}^{(k)}\}_{k=1}^{K}$, screening studies with small $\|\boldsymbol{\delta}^{(k)}\|_{1}$ is relatively simple. We can directly calculate each $\|\widetilde{\boldsymbol{\delta}}^{(k)}\|_{1}$ and set a threshold to reject the dissimilar studies. As implied by Corollary 1, $h_{1}$ should be no more than the single-task error [zhang2022transfer, huang2023estimation], here we set the threshold as $t_{1}\sqrt{\log p/n_{0}}$, where $t_{1}$ is a thresholding parameter. The first screening set is then defined as $\widehat{\mathcal{C}}_{1}=\{k:\|\widetilde{\boldsymbol{\delta}}^{(k)}\|_{1}\leq t_{1}\sqrt{\log p/n_{0}}\}$. We will show that under some regularity conditions, appropriate choice of $t_{1}$ will lead to consistent selection conforming with the requirement $\|\boldsymbol{\delta}^{(k)}\|_{1}\leq h_{1}$.
@@
The second requirement of the oracle transferable set $\mathcal{C}_{h}$ is to reject source studies with tiny information $n_{k}\mathbb{E}[f_{i}^{(k)}]$ compared to the target $n_{0}\mathbb{E}[f_{i}^{(0)}]$ around the $\tau$-th quantile. We employ a kernel-based approximation using the Powell bandwidth $b_{k}$ [dai2021inference, huang2023estimation, koenker2005quantile] to estimate each $\mathbb{E}[f_{i}^{(k)}]$, with the choice of $b_{k}$ detailed in the supplementary materials. Specifically, for each study we calculate
@@
$\widetilde{f}^{(k)}(\widetilde{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)})=\frac{1}{2b_{k}n_{k}}\sum_{i=1}^{n_{k}}\mathbb{I}\left(|y_{i}^{(k)}-(\boldsymbol{x}_{i}^{(k)})^{\top}(\widetilde{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)})|\leq b_{k}\right).$ (7)
[…]
Note that the calculation of estimators $\widetilde{\boldsymbol{\beta}},\{\widetilde{\boldsymbol{\delta}}^{(k)}\}_{k=1}^{K}$ is exactly the first round of iterative procedure for solving the $\ell_{1}$-minimization transfer framework (3). Hence here, we propose a unified one-shot algorithm for transferable set detection and parameter prediction, which is summarized in Algorithm 1. Note that for complex practical scenarios, such as strong distribution shift or heavy-tailed distributions, the procedure need to be iterated multiple times until the detected set has stabilized.
@@
Input: Target data $\{((\boldsymbol{x}_{i}^{(0)})^{\top},y_{i}^{(0)})_{i=1}^{n_{0}}\}$, source data $\{((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)})_{i=1}^{n_{k}}\}_{k=1}^{K}$, thresholds $t_{1}$ and $t_{2}$, tuning parameters $\lambda_{\boldsymbol{\beta}}$ and $\{\lambda_{k}\}_{0\leq k\leq K}$, bandwidths $\{b_{k}\}_{0\leq k\leq K}$.
@@
    Compute $\widetilde{\boldsymbol{\delta}}^{(k)}\leftarrow\ell_{1}\text{-QR }\text{ on }[((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)}-(\boldsymbol{x}_{i}^{(k)})^{\top}\widetilde{\boldsymbol{\beta}})_{i=1}^{n_{k}},\lambda_{k}],1\leq k\leq K$.
@@
    Compute $\{\widetilde{f}^{(k)}(\widetilde{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)})\}_{0\leq k\leq K}$ with bandwidths $\{b_{k}\}_{0\leq k\leq K}$ via (7).
@@
[…]
  5. 5.
    Determine the transferable set $\widehat{\mathcal{C}}=\widehat{\mathcal{C}}_{1}\cap\widehat{\mathcal{C}}_{2}$, then solve the pooled $\ell_{1}$-QR
@@
We next establish the detection consistency of $\widehat{\mathcal{C}}$ towards the oracle but unknown transferable set $\mathcal{C}_{h}$. To ensure the identifiability of source studies satisfying the oracle requirement in (2), we need to assume a significant gap of quantile coefficients and $\tau$-th densities, where similar conditions are also assumed in [tian2022transfer, zhang2022transfer, huang2023estimation].
@@
For each source study $k$, let $\Omega_{1}^{(k)}:=C_{3}\frac{s_{k}}{\mathbb{E}[f_{i}^{(k)}]}\sqrt{\frac{\log p}{n_{k}\wedge n_{0}}}+2\eta_{k}$, where $C_{3}$ is a positive constant specified in the supplementary materials. Denote the set $\mathcal{C}_{h_{1}}:=\{k\in\{1,\dots,K\}:\|\boldsymbol{\delta}^{(k)}\|\leq h_{1}\}$. $\mathcal{C}_{h_{1}}$ satisfies that
@@
where $r_{k}$ is defined in Theorem 1 ($r_{0}=0$), and $C_{4}$ is a positive constant specified in the supplementary materials. Denote the set $\mathcal{C}_{h_{2}}:=\{k\in\{1,\dots,K\}:(n_{0}\mathbb{E}[f_{i}^{(0)}])/(n_{k}\mathbb{E}[f_{i}^{(k)}])\leq h_{2}\}$. $\mathcal{C}_{h_{2}}$ satisfies that
$\inf_{k\in\mathcal{C}_{h_{2}}}\frac{n_{k}(\mathbb{E}[f_{i}^{(k)}]-\Omega_{2}^{(k)})}{n_{0}(\mathbb{E}[f_{i}^{(0)}]+\Omega_{2}^{(0)})}>\sup_{k\in\mathcal{C}_{h_{2}}^{c}}\frac{n_{k}(\mathbb{E}[f_{i}^{(k)}]+\Omega_{2}^{(k)})}{n_{0}(\mathbb{E}[f_{i}^{(0)}]-\Omega_{2}^{(0)})}.$
@@
Similar to previous transfer literature [tian2022transfer, zhang2022transfer], Conditions 5 and 6 assume the parameter contrast $\|\boldsymbol{\delta}^{(k)}\|_{1}$ and the population $\tau$-th density $\mathbb{E}[f_{i}^{(k)}]$ exhibit sufficient theoretical separation between transferable and non-transferable source studies, which exactly satisfies the requirement of the oracle transferable set (2). To be more specific, the value $\Omega_{1}^{(k)}$ is the upper error bound of $\|\widehat{\boldsymbol{\delta}}^{(k)}-\boldsymbol{\delta}^{(k)}\|_{1}$, while the value $\Omega_{2}^{(k)}$ is the upper error bound of $|\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\widetilde{f}_{i}^{(k)}(\widetilde{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)})-\mathbb{E}[f_{i}^{(k)}]|$.
@@
Assuming Conditions 1-6 and assumptions in Theorem 1 for all source studies, choose the thresholding parameters $t_{1},t_{2}$ in Algorithm 1 as
@@
$\inf_{k\in\mathcal{C}_{h_{2}}}\frac{n_{k}(\mathbb{E}[f_{i}^{(k)}]-\Omega_{2}^{(k)})}{n_{0}(\mathbb{E}[f_{i}^{(0)}]+\Omega_{2}^{(0)})}$ $\geq t_{2}\geq\sup_{k\in\mathcal{C}_{h_{2}}^{c}}\frac{n_{k}(\mathbb{E}[f_{i}^{(k)}]+\Omega_{2}^{(k)})}{n_{0}(\mathbb{E}[f_{i}^{(0)}]-\Omega_{2}^{(0)})}.$
Then for the detected transferable set $\widehat{\mathcal{C}}$ in Algorithm 1, for any $\varepsilon>0$, there exists a positive constant $N(\varepsilon)$ such that when $\min_{k}n_{k}>N(\varepsilon)$, $\mathbb{P}(\widehat{\mathcal{C}}=\mathcal{C}_{h})\geq 1-\varepsilon$. Consequently, our practical $\widehat{\boldsymbol{\beta}}$ obtained from Algorithm 1 can enjoy the same $\ell_{1}/\ell_{2}$-convergence rates as the oracle estimator $\widehat{\boldsymbol{\beta}}_{ora}$ in Corollary 1.
@@ 4.1 Debiased Estimator based on Neyman Orthogonality
We now turn to the statistical inference for the target parameter, $\boldsymbol{\beta^{*}}$, focusing specifically on the first component, $\beta^{*}_{1}$, without loss of generality. Previous studies [tian2022transfer, li2023estimation, huang2023estimation] addressed this inference task building on the principles of debiased lasso [zhang2014confidence, van2014asymptotically]. Unfortunately, these methods only achieve $\sqrt{n_{0}}$-normality for the debiased estimators, offering little improvement over debiasing based solely on the target study. The key reason for this limitation is that their debiasing procedures need the covariance matrix to be identical for all samples, a condition that is violated when covariate shift occurs. Also, they do not offer any measures to adjust for the parameter contrast $\boldsymbol{\delta}^{(k)}$. The $\sqrt{n_{0}}$-normality is suboptimal, especially given the wealth of transferable information from source studies. The challenge, then, is how to effectively leverage the source information to enhance inference efficiency, particularly under distributional shift.
To address this challenge, we propose a novel approach motivated by the Neyman orthogonal idea [belloni2019valid]. Denote $\boldsymbol{v}_{-1}$ as the last (p1)(p-1)-dimensions of the vector $\boldsymbol{v}\in\mathbb{R}^{p}$. We can rewrite the original model (1) as a partially linear quantile regression
@@
The high-dimensional nuisance parameter $\boldsymbol{w}_{-1}^{(k)}$ will bring trouble for identifying $\beta^{*}_{1}$, since the first-order derivative of (8) with respect to this nuisance parameter at $\alpha=\beta^{*}_{1}$ is non-zero. If conducting valid inference on $\beta^{*}_{1}$ directly with moment condition (8), the estimator for $\boldsymbol{w}_{-1}^{(k)}$ needs to converge faster than $1/\sqrt{n_{k}}$, which is difficult to achieve. To address this problem, we adopt the Neyman orthogonality to make the moment condition immune to the first-order error from the nuisance estimator. The technique is based on the linear projection of the regressor of interest $x_{i,1}^{(k)}$ on the $\boldsymbol{x}^{(k)}_{i,-1}$ weighted by the $\tau$-th density $f_{i}^{(k)}$:
@@
$\psi_{i}^{(k)}(\alpha,\boldsymbol{w}_{-1}^{(k)},\boldsymbol{\theta}_{0}^{(k)},f_{i}^{(k)}):=\left[\mathbb{I}\left\{y_{i}^{(k)}\leq x_{i,1}^{(k)}\alpha+(\boldsymbol{x}_{i,-1}^{(k)})^{\top}\boldsymbol{w}_{-1}^{(k)}+r_{i}^{(k)}\right\}-\tau\right]v_{0i}^{(k)},$ (10)
@@
and solve for $\mathbb{E}\sum_{i=1}^{n_{k}}[\psi_{i}^{(k)}(\alpha,\boldsymbol{w}_{-1}^{(k)},\boldsymbol{\theta}_{0}^{(k)},f_{i}^{(k)})]=0$, where this moment condition can satisfy the orthogonality with respect to high-dimensional nuisance parameters $\boldsymbol{w}_{-1}^{(k)}$ and $\boldsymbol{\theta}_{0}^{(k)}$, i.e.
$\left.\partial_{\boldsymbol{w}}\mathbb{E}\sum_{i=1}^{n_{k}}\left[\psi_{i}^{(k)}(\beta_{1}^{*},\boldsymbol{w},\boldsymbol{\theta}_{0}^{(k)},f_{i}^{(k)})\right]\right|_{\boldsymbol{w}=\boldsymbol{w}_{-1}^{(k)}}=\left.\partial_{\boldsymbol{\theta}}\mathbb{E}\sum_{i=1}^{n_{k}}\left[\psi_{i}^{(k)}(\beta_{1}^{*},\boldsymbol{w}_{-1}^{(k)},\boldsymbol{\theta},f_{i}^{(k)})\right]\right|_{\boldsymbol{\theta}=\boldsymbol{\theta}_{0}^{(k)}}=0.$
Consequently, the estimators for $\boldsymbol{w}_{-1}^{(k)}$ and $\boldsymbol{\theta}_{0}^{(k)}$ will only introduce a second-order bias, allowing the use of high-dimensional sparse estimates. Note that this approach can avoid the computation of precision matrices and free the identical covariance matrix assumption across studies, which exactly accommodates the circumstance under distribution shift.
@@
We outline several points here for practical implementation. Firstly, we estimate the $\tau$-th density $f_{i}^{(k)}$ by $\widehat{f}_{i}^{(k)}=2h^{(k)}/[(\boldsymbol{x}_{i}^{(k)})^{\top}\widehat{\boldsymbol{\beta}}^{(k)}_{\tau+h^{(k)}}-(\boldsymbol{x}_{i}^{(k)})^{\top}\widehat{\boldsymbol{\beta}}^{(k)}_{\tau-h^{(k)}}]$, where $\widehat{\boldsymbol{\beta}}^{(k)}_{\tau\pm h^{(k)}}$ are computed via $\ell_{1}$-QR at the $(\tau\pm h^{(k)})$-th quantiles with tuning parameters $\lambda_{\tau}^{(k)}$ on the $k$-th study respectively. The bandwidth parameter, $h^{(k)}$, is set to $\min\{n_{k}^{-1/6},\tau(1-\tau)/2\}$, as suggested by [koenker2005quantile, giessing2023debiased]. Next, each high-dimensional projection parameter $\boldsymbol{\theta}_{0}^{(k)}$ is determined by $\ell_{1}$-penalized weighted least squares with tuning parameters $\lambda_{\boldsymbol{\theta}}^{(k)}$
$\widehat{\boldsymbol{\theta}}^{(k)}\in\arg\min_{\boldsymbol{\theta}}\sum_{i=1}^{n_{k}}\left[\widehat{f}^{(k)}_{i}x_{i,1}^{(k)}-\widehat{f}^{(k)}_{i}(\boldsymbol{x}^{(k)}_{i,-1})^{\top}\boldsymbol{\theta}\right]^{2}+\lambda_{\boldsymbol{\theta}}^{(k)}\|\boldsymbol{\theta}\|_{1}.$ (11)